Using charpits method


(p+q)(z-px-qy)=1


1
Expert's answer
2021-01-14T16:13:30-0500

"(p+q)(z-px-qy) =1\\
f=(p+q)(z-px-qy) -1=0"

Charpit's auxillairy equation is given as;

"\frac{dx}{-\frac{\partial f}{\partial p}}=\frac{dy}{-\frac{\partial f}{\partial q}}=\frac{dz}{-p\frac{\partial f}{\partial p}-q\frac{\partial f}{\partial p}}=\frac{dp}{\frac{\partial f}{\partial x}+p\frac{\partial f}{\partial z}}=\frac{dq}{\frac{\partial f}{\partial y}+q\frac{\partial f}{\partial z}}\\
\frac{\partial f}{\partial x}=-p^2-pq\\
\frac{\partial f}{\partial y}=-pq-q^2\\
\frac{\partial f}{\partial z}=p+q\\
\frac{\partial f}{\partial p}=z-xq-yq-2xp\\
\frac{\partial f}{\partial q}=z-xp-yp-2yq"

Now, the Charpit's equation is;

"\frac{dp}{0}=\frac{dq}{0}=\frac{dz}{2x(p^2+pq)+2y(q^2+pq)-2zp}=\frac{dy}{2q+xp+yp-z}=\frac{dx}{2xp+yq+xq-z}\\
\text{The first fraction implies, }\\
dp=0 \implies p=a\text{ By integrating }\\
\text{Similarly, the second fraction implies }\\
dq=0 \implies q=b\text{ By integrating }\\
\text{Now substitute into }dz=pdx+qdy\\
dz=adx+bdy\\
\text{Integrate through }\\
z=ax+by+c\\
\text{Substitute the value of z,p and q into f to get c}\\
(a+b)(ax+by+c-ax-by)=1\\
c=\frac{1}{a+b}\\
\implies z=ax+by+\frac{1}{a+b}\\
\implies (a+b)(z-ax-by) =1"